OpenStudy (anonymous):

Problem 14.4 A Prism A rectangular prism's surface area is equal to 324 cm2. If its length and width are 12 cm and 8 cm respectively, find the prism's height. Question 3 Determine the maximum number of 20 cm by 15 cm Christmas cards you can make from a 3 m by 2 m cardboard sheet, if each card has four pages (i.e. 2 leaves). Question 4 A car's engine has 6 cylinders, each of diameter 9.6 cm and height 10.4 cm. Find the capacity of the engine's cylinders in cc (cubic centimetres).

OpenStudy (anonymous):

1. 2*12*8+2*8*h+2*12*h=324cm^2 solve and find h?

OpenStudy (anonymous):

2. no of cards = 6/(0.2*0.15)

OpenStudy (anonymous):

3. capacity= pi*(9.6/2)^2*10.4*6 cc

OpenStudy (anonymous):

Thanks so much!